OpenStudy (anonymous):

a particle is thrown upwards.it's velocity at half of the height is 10m/s,then calculate the maximum hieght attained by it

OpenStudy (matt101):

We know acceleration = -9.8 m/s^2 and final velocity = 0 m/s. If we treat 10 m/s as the initial velocity, the total distance it travels until its velocity reaches 0 is HALF the maximum height, since it is traveling at 10 m/s HALFWAY through its trip. Use vf^2 = vi^2 + 2ad: 0 = 10^2 - 2(9.8)(0.5hmax) hmax = 10.2 m

OpenStudy (anonymous):

my answer is also same as above one, i..e., max. height attained= 10.2m

OpenStudy (anonymous):

listen.... i am a 9th grade and cannot understand your equations can someone give me a answer which i can understand ....plz.......i have to submit my homework....

OpenStudy (foolaroundmath):

do you know the three equations if motion for constant acceleration? i.e. v = u + at s = ut + 0.5at^2 v^2 - u^2 = 2as

OpenStudy (anonymous):

no.....i don't just explain plz.

OpenStudy (anonymous):

whats s,u,a,t are they short forms

OpenStudy (foolaroundmath):

u = intial velocity, v= final velocity a = acceleration t=time s = displacement

OpenStudy (anonymous):

and what do you mean by this thing ^2

OpenStudy (foolaroundmath):

its the exponent .. x^y = x raised to y

OpenStudy (anonymous):

so it means square of something

OpenStudy (matt101):

Do you understand how we got the answer? I'll rewrite my answer using the equation tool to make it clearer if you need and provide more explanation.